OpenStudy (anonymous):

radical 24 divided by radical 6

OpenStudy (teddyiswatshecallsme):

wouldn't that just be a radical 4? or r = 4? I dont know this math so idk....

OpenStudy (anonymous):

Have you tried multiplying the top and bottom by the 'conjugate'

OpenStudy (anonymous):

Scratch that, use \[\frac{ \sqrt{24} }{ \sqrt{6} } \implies \sqrt{\frac{ 24 }{ 6 }}\]

OpenStudy (anonymous):

That should give you the right answer :)

OpenStudy (anonymous):

\[\sqrt{\frac{ 24 }{ 6 }} \implies \sqrt{4} \implies 2\]

OpenStudy (anonymous):

nice to now have the same problem